Previous

Calendar

August

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

29

JD210

30

JD211
Site Visit (JB,DK,SL,NT): Details

SL and NT started diurnal soil respiration measurements at dawn; JB and DK helped

at HYDRA: at 9:45 MST, dropped dry ice into HYDRA box (JB)

Trade in jeep and pick up suburban

31

JD212

1

JD213
Site Visit (JB, RKM, SpB, RaM, RM): Details , Photos (9)

JB replaced Jia's electrical board at west 90

at ~12:00 MST,cleaned krypton hygrometer optics

downloaded from soil, Tcs, RAD, PROF, MET, NCAR Tcs

from 12:15-12:30 MST, adjusted diffuser so the par sensor was shaded

WSU leaves (to return next week)

2

JD214
Site Visit (JB,SL): Details

Brought 3 N2 cylinders to the trailer (CC125575, CC193045, CC198701)

Added salt to both grounding pipes

closed main valves on both hydra gas cylinders, and powered down hydra

swapped 15 flasks

3

JD215

4

JD216
Note: high-freq noise in the 21.5m csat (see between 14-15 MDT)??

5

JD217

6

JD218
Site Visit (JB, SpB): Details , Photos (17)

at 11:11 MST, restarted HYDRA: 16 inlets connected (could not find "U-Como1" so B4 = short tube near box). C1/C3 are not working. Ref=zero air, C2=349 ppm, C4=384 ppm

at 14:43 MST, slightly increased the flow rate of the HYDRA ref gas (zero air)

Adjusted position of VB-C10R webcam. Tested "forestcam.colorado.edu".

7

JD219
SpB Off (appendectomy)

8

JD220
SpB Sick Day
at 16:21 MST, hi-rate data archiving stopped on russter2.

9

JD221
Site Visit (JB,SL,SLS,DK,Ramesh,LL): Details

at 9:10 MST, hi-rate data archiving re-started on russter2.

JB added more water and salt to both grounding tubes

JB adjusted diffuser (see details)

JB swapped 15 flasks

JB replaced mag. perch. trap and indicator tube for the 6251 (see details)

JB checked cylinder pressures @ hydra (see details)

new ozone target (AT)

10

JD222

11

JD223
WSU sf6 (PM)

12

JD224
WSU sf6 (PM)

13

JD225
WSU sf6 (PM)

at 16:00 MDT, SpB to Dr.

14

JD226
SpB Sick Day
WSU sf6 (PM)

15

JD227
SpB Sick Day
WSU packs up (10 am - whenever)

16

JD228
Site Visit (JB,SL): Details

JB downloaded from RAD, PROF, MET, NCARTC, UCB, UCAN, Tcs, Soil

@ 10:00 MST JB adjusted the shading band to ~27.52 degrees (this is the solar noon sun angle for 8/19, which was the closest day for which I had sun info

JB swapped 15 flasks

see details for cylinder pressure info

SpB Sick Day

17

JD229
JB has doctor's appointment

18

JD230

19

JD231

20

JD232
Site Visit (JB,SL,MP,RKM,LL,Steve Semmer,John Militzer): Details

JB, RKM, MP, SS, JM deployed the photodiode "wand" sensors in MP's spruce and fir trees

JB did a "collect all" on the 23xsoil logger

21

JD233
at 14:15 MDT, SpB to Dr.

22

JD234

23

JD235
Site Visit (JB,JH,MP,SpB,SLS): Details

at ~11:30 MST, rotated shadow band by ~5 deg CLOCKWISE

JB swapped 15 flasks

at 9:20 MST, ATI 2-D sonic added to channel 216 serial port on the back of the quacker (start ops44)

from 12-12:30 MST at HYDRA: changed C2 from 349ppm (JA02103) to 417ppm (JJ16580).

see details for cylinder pressures

at ~9:20 MST, security update to russter2 (rebooted).

24

JD236

25

JD237

26

JD238

27

JD239 (cool, overcast and rainy)
Site Visit (JB): Details

brought 4th N2 cylinder up to the trailer from N324

removed the TRAM trolley from its track for SO

got the Vaisala HMP35AC T/RH sensor (s/n S0630018) from the north canopy tower for the new aspiration unit

checked connections on the 21.5m C-SAT and into the FAST logger, couldn't see anything out of place...call Campbell

see details for newest cylinder pressure info

see details for HMP35AC 23x wiring scheme

28

JD240 (windy,cool,afternoon rain)
SpB Off
Site Visit (JB,JH,MP,NT): Details

POWER OUTAGE TO TOWER AND TRAILER FROM ~9:53-9:55 MDT

downloaded from RAD, PROF, MET, NCARTcs, SOIL, and Tcs

29

JD241
mailed broken AM25T back to Campbell (RMA#5257)

got new valve for Hewlett Packard N2 regulator #8507-0407, started new leak check (this valve was pro bono publico from AirGas)

30

JD242 (cool, heavy rain)
Site Visit (JB,NT): Details

Xcel Energy planned power outage at MRS from ~3:00-3:30 MDST

JB swapped 15 flasks in a heavy downpour with thunder and lightning

31

JD243
at 17:54 MDT: ndaq process dies, HYDRA data stop being collected.

JB Off

1

JD244